# Excitation Waveforms for Maximum Instantaneous Power Delivery

### v1

@@ II Historical Background
Supplementary to [7] and in alignment with [8], optimal excitations for dipole arrays aimed at maximizing the amplitude of the transient radiated electric field at a specified time and position in the far field are detailed in [9], and maximizing energy density in a given time interval at a fixed far-field position are detailed in [2]. As with [7], the MoM is employed to calculate the currents on all dipoles, taking into account the effects of mutual coupling. Piecewise sinusoidal expansion modes are used, reflecting the approach in [7]. The relationship between the total radiated field and terminal voltages is established by superposing the far fields of each terminal. This analysis uses a thin-wire approximation for the dipoles and delta gaps to model their feed-voltage sources. The degrees of freedom in this study encompass the magnitudes and phases of the terminal voltages. Furthermore, the investigation is extended to the synthesis of optimal sidelobe patterns.
@@
Moving to communication systems, optimizing the transient waveform should consider both the transmitting and receiving antennas, in contrast to the work mentioned above. In ultra-wideband radio systems, the crucial metric is the maximum amplitude of the received antenna voltage, its sharpness, or energy, as addressed in [35]. Unbeatable target metrics are achieved by finding the best-performing waveform considering the constraints on energy in different stages of the channel. General optimization results are derived for arbitrary antennas, considering the effects of the generator and load impedances using variational methods. The approach is demonstrated with dipoles using MoM [^3] Piecewise sinusoidal basis is used as well, as in [7, 9]. solutions providing the necessary transfer functions and input impedances from the EFIE analysis with the study enriched by closed-form results for short-dipole antennas. Ideally, an ultra-wideband transmitter’s radiated power spectral density should be as flat as possible. However, in reality, ultra-wideband systems are limited by the performance of the antennas which is far from ideal [35].
@@ III Framework
This section develops and verifies the computational approach for optimal excitation and showcases it on pulsed radiation in a chosen far-field direction and time. Various goals and objective functions with multiple constraints can be considered within the approach presented in this article. The practical usage of the method is given in Section IV where several applications are considered.
@@ III-A Methodology
    First, an antenna with defined ports and feeding transmission lines has to be chosen. The setup is analyzed in an arbitrary EM (EM) solver, which can compute the far field in the desired direction for a given frequency and incident waves on the feed lines to the antenna ports. The EM solver provides a set of far-field values. An arbitrary method, capable of evaluating target quantities for a given excitation, can be used, i.e., finite differences in time domain (FDTD) in CST studio suite [37] (CST).
@@
    $\boldsymbol{E}_{\infty}(\boldsymbol{r},\omega)=\mathbf{H}(\boldsymbol{r},\omega)\,\mathbf{a}(\omega)\mathrm{e}^{-\mathrm{j}\omega\dfrac{r}{c}},$ (3)
    where $\mathbf{H}$ is a far-field transfer matrix obtained from the chosen solver. An example of getting the transfer matrix is the combination of the EFIE and the MoM detailed in Appendix A. Apart from the quantities needed for optimization, antenna properties such as the reflection coefficient or gain at given frequencies can be examined to provide insight into the spectral selectivity of the antenna.
@@
    The spectral functions representing the necessary operators for optimization, i.e., total incident energy, and the far field are projected onto a set of continuous-band-limited basis functions. The set of basis functions $\xi_{n}(\omega)$ is described in Appendix B. The projection results in expansion coefficients of the basis function grouped in column vector $\mathbf{q}$, where
@@
    The vector is constructed in segments with each segment comprising the expansion coefficients of the basis functions for each feed line $p$. Related matrices $\mathbf{W}$ (incident energy) and $\mathbf{S}$ (far-field power) are derived from (4) and (2) together with (3), respectively. See Appendix B for details.
@@
    The optimization task is resolved using standard methods of the QCQP. For more details, see [38].
@@ III-B Verification
The steps mentioned above are applied to an example of a single dipole antenna. The calculation is simultaneously performed in the Antenna Toolbox for MATLAB [39] (AToM) and CST. The dipole with a length of $15\,\mathrm{mm}$ and a width of $3\,\mathrm{mm}$ is oriented along the $z$ axis. The port is placed in the middle of the dipole and is connected to an ideal $50\,\mathrm{\Omega}$ feed line. Radiation in the spherical coordinate system is studied in the direction $\varphi=0$ and $\theta=\pi/2$.
The antenna including the port is given and all steps visualized in Fig. 1 can be performed. Choosing AToM and CST as the EM solvers and computing the far field for unity excitation at all frequencies results in Fig. 2.
@@ IV-A1 Singleport Antennas
In [42], the conical spiral antenna is proposed for ground-penetrating radars, an ultra-wideband application. The conical spiral radiates a circularly polarized field compared to the other designs which operate in a linearly polarized regime. The antenna proposed in [42] is scaled to fit the circumscribing sphere mentioned above. The simulation is again performed in CST and shown in Fig. 7.
@@
The periodic logarithmic antenna, recognized as an ultra-wideband antenna [43], is extensively utilized in electromagnetic emission and immunity testing [6]. Therefore, a wire dipole array is a potential candidate for pulsed radiation and is compared with other antennas. The array is designed to fit the circumscribing sphere and exhibits a decay of $0.9$ in element lengths and an opening angle of $15$ degrees. The strip width equals $0.017$ times the length of the longest element and the number of elements is 20. The antenna is simulated using AToM and shown in Fig. 8.
@@
The comparison of the antennas and the finding of the optimal excitation using the approach proposed in this paper show that antenna design and pulse generation capabilities are closely related. The information that can be provided by this computational approach can provide essential information in designing antennas that perform well in pulsed radiation.
@@ IV-A2 Multiport Systems
The extension to multiport antennas or arrays is straightforward, as with dipole arrays within this section. The matrix formulation is also shown in Appendices A and B.
@@
Considering the same input energy, the end-fire array increases the performance by more than a factor of four as compared to a single dipole. In the case of a broadside array, the performance boost reaches a factor of six. The difference may be due to the coupling of the field with the elements of the arrays which can be deduced from the optimal excitation waveforms in Fig. 13.
@@ IV-B Switching Antiferromagnetic Memory
Figure 15: Incident plane waves at the position of the observation: (o) Pulse used in [44]. (a) Optimal excitation for the maximum peak. (b) Optimal excitation for the maximum peak with a constraint on the incident wave to concentrate $90\,\%$ of energy in the $2~{}\mathrm{ps}$ interval. (c) Optimal excitation for the maximum peak with a constraint in the target field to concentrate $90\,\%$ of energy in the $14~{}\mathrm{ps}$ interval. (d) Optimal excitation for maximum peak with both constraints in (b) and (c).
@@
Figure 16: Magnitude of electric field in CuMnAs for excitation pulses (o)–(d) in Fig. 15.
@@
One of the optimal pulses presented corresponds to the maximum magnitude of the electric field in CuMnAs in $t=0$ which is denoted as (a). The optimization problem is identical to (6) other than matrix $\mathbf{S}$ being replaced by $\mathbf{E}^{\mathrm{H}}\mathbf{E}$, where the electric field at the target point is represented as $\boldsymbol{E}_{\mathrm{CuMnAs}}(\omega)=\mathbf{E}(\omega)\mathbf{q}$ in the frequency domain, and $\mathbf{q}$ corresponds to the expansion of incident plan wave $a(\omega)$ into the basis functions of Appendix B, as done in (5). The relation is computed using CST which allows us to compute the electric field at a given point as a function of frequency considering unity excitation. Mapping the set of the given basis functions results in the desired matrix. The matrix representing the total energy of the incident plane wave is derived from (4) as demonstrated in Appendix B, with the only variation being the inclusion of the free space impedance in the denominator.
@@
where $T=1\,\mathrm{ps}$ is the half-time interval and the asterisk represents a complex conjugation. In the frequency domain, the sinc function represents the process of time windowing. The matrix obtained is the outcome of the procedure illustrated in Appendix B regarding total input power. The corresponding QCQP reads
@@
The combination of both constraints for excitation in (b) and for the target field in (c) results in another optimization problem whose solution is presented in (d). Since optimization has more constraints than the previous solutions (a), (b), and (c), performance is lower, but still higher than that of the original pulse (o). The excitation mostly spans the narrow interval and the electric field in CuMnAs is also mainly accumulated in the given interval. The disparity between (o) and (d) is minimal, suggesting that enhancement can only be achieved through topological changes.
@@ V Conclusion
The tool can serve and provide quantitative results in many scenarios, including pulsed radiation or antiferromagnetic memory switching with pulse lasers, as discussed in this paper, as well as the creation of hotspots in lossy materials using antenna arrays for applications such as hyperthermia therapy in medicine.
@@
When the optimized pulsed response is unsatisfactory, the only way to further improve the performance is topology optimization of the radiating body. Topology optimization would change the design according to the results computed by the proposed tool, though such a procedure would be computationally intensive. However, this would open the way to the automated design of pulsed radiators, antiferromagnetic memories, or other devices that rely on pulses in the time domain.
@@ Appendix A Matrix Operators from EFIE and MoM
A representative example of an EM solver capable of providing required antenna characteristics is AToM based on EFIE and MoM. The matrix representation of multiple quantities is presented in [38, Appendix A]. The extraction of essential metrics to the manuscript is presented in this appendix. Multiple ports are considered.
@@
$\boldsymbol{E}_{\infty}(\boldsymbol{r},\omega)=\mathbf{F}(\boldsymbol{r},\omega)\mathbf{I}(\omega)\dfrac{\mathrm{e}^{-\mathrm{j}\dfrac{\omega}{c}r}}{r},$ (11)

### v2

# Optimization of Excitation Waveforms for Maximum Instantaneous Field Intensity

@@ I Introduction
An essential advantage of the proposed method lies in its ability to assess whether a given system configuration can achieve the desired pulsed performance. The method provides critical insights into whether optimization efforts can be confined to pulse shaping or whether a redesign of the system geometry is necessary. This dual capability offers valuable guidance for both waveform optimization and structural design, facilitating the decision-making process.
While co-design approaches that simultaneously optimize waveform and antenna design may approach globally optimal solutions, this work focuses specifically on waveform optimization for fixed antenna designs. By decoupling waveform optimization from antenna design, the method ensures broad applicability and low computational complexity.
@@
Although this work primarily focuses on optimizing excitation waveforms to achieve power metrics in far-field scenarios. The proposed framework is easily adjustable for near-field energy focusing or similar applications where precise waveform shaping is crucial, given the knowledge of the impulse response in the respective region. However, this work does not address the practical aspects of designing pulsed systems but instead provides a versatile theoretical methodology that is applicable in diverse contexts.
@@ II Historical Background
Supplementary to [7] and in alignment with [8], optimal excitations for dipole arrays aimed at maximizing the amplitude of the transient radiated electric field at a specified time and position in the far field are detailed in [9], and maximizing energy density in a given time interval at a fixed far-field position are detailed in [2]. As with [7], the MoM is employed to calculate the currents on all dipoles, taking into account the effects of mutual coupling. Piecewise sinusoidal expansion modes are used, reflecting the approach in [7]. The relationship between the total radiated field and terminal voltages is established by superposing the far fields of each terminal. This analysis uses a thin-wire approximation for the dipoles and delta gaps to model their feed-voltage sources. The dof in this study encompass the magnitudes and phases of the terminal voltages. Furthermore, the investigation is extended to the synthesis of optimal sidelobe patterns.
@@
Moving to communication systems, optimizing the transient waveform should consider both the transmitting and receiving antennas, in contrast to the work mentioned above. In ultra-wideband (UWB) radio systems, the crucial metric is the maximum amplitude of the received antenna voltage, its sharpness, or energy, as addressed in [35]. Unbeatable target metrics are achieved by finding the best-performing waveform considering the constraints on energy in different stages of the channel. General optimization results are derived for arbitrary antennas, considering the effects of the generator and load impedances using variational methods. The approach is demonstrated with dipoles using MoM [^3] Piecewise sinusoidal basis is used as well, as in [7, 9]. solutions providing the necessary transfer functions and input impedances from the EFIE analysis with the study enriched by closed-form results for short-dipole antennas. Ideally, an UWB transmitter’s radiated power spectral density should be as flat as possible. However, in reality, UWB systems are limited by the performance of the antennas which is far from ideal [35].
@@ III Framework
This section develops and verifies the general computational approach for optimal excitation and showcases it on particular example of pulsed radiation in a chosen far-field direction and time. Various goals and objective functions with multiple constraints can be considered within the approach presented in this article. The essential user input for the computation is the relation between the excitation and the target metric, i.e., a solution to Maxwell’s equations. Knowing the system response, almost arbitrary linear scenarios can be studied, including optimization of radiated field intensity, power delivery efficiency, or waveform shaping tailored to specific spatial or temporal requirements. This versatility makes the framework applicable to scenarios such as pulse radiation enhancement and electromagnetic field focusing in localized regions. In addition to the examples given in the manuscript, notable applications might be the wireless power delivery from the port of an antenna to the port of another antenna and the creation of hot spots in the near field. The practical usage of the method is given in Section IV where several applications are considered.
@@ III-A Methodology
    First, an antenna with defined ports and feeding transmission lines has to be chosen. The setup is analyzed in an arbitrary EM (EM) solver, which can compute the far field in the desired direction for a given frequency and incident waves on the feed lines to the antenna ports. The EM solver provides a set of far-field values. An arbitrary method, capable of evaluating target quantities for a given excitation, can be used, i.e., finite differences in time domain (FDTD) in CST studio suite [37] (CST). This part of the process is computationally the most time-consuming. The computational demands grow with wider frequency ranges $f\in\left[f_{\mathrm{min}},f_{\mathrm{max}}\right]$, but these, on the other hand, introduce more dof in the waveform design providing greater flexibility in shaping the radiated pulses.
@@
    $\boldsymbol{E}_{\infty}(\boldsymbol{r},\omega)=\mathbf{H}(\boldsymbol{r},\omega)\,\mathbf{a}(\omega)\mathrm{e}^{-\mathrm{j}\omega r/c},$ (3)
    where $\mathbf{H}$ is a far-field transfer matrix obtained from the chosen solver and $a(\omega)$ represents incident waves in the frequency domain (for example, power waves in the feeding transmission lines), i.e., Fourier transform of $a(t)$. An example of getting the transfer matrix is the combination of the EFIE and the MoM detailed in Appendix A. Apart from the quantities needed for optimization, antenna properties such as the reflection coefficient or gain at given frequencies can be examined to provide insight into the spectral selectivity of the antenna.
@@
    The spectral functions representing the necessary operators for optimization, i.e., total incident energy, and the far field are projected onto a set of continuous-band-limited basis functions. The projection follows Galerkin method [38, 13]. The set of basis functions $\xi_{n}(\omega)$ is described in Appendix B. For all considered examples, number of basis functions [^4] Later, the number of basis functions is the number of dof in the optimization and determines the time needed for construction $\mathcal{O}(N^{2})$ of matrices and solution $\mathcal{O}(N^{3})$. $N\leq 120$ was sufficient. The projection results in expansion coefficients of the basis function grouped in column vector $\mathbf{q}$, where
@@
    The vector is constructed in segments with each segment comprising the expansion coefficients of the basis functions for each feed line $p$. Related matrices $\mathbf{W}$
    $W_{\mathrm{inc}}=\mathbf{q}^{\mathrm{H}}\mathbf{W}\mathbf{q}$ (6)
    (incident energy) and $\mathbf{S}$ (far-field power) are derived via Galerkin method from (4) and (2) together with (3), respectively.
    As an example, the incident energy (4) is represented by matrix $\mathbf{W}$, the components of which read
    $\mathbf{W}_{pm,qn}=\dfrac{\delta_{pq}}{4\pi}\int\limits_{-\infty}^{\infty}\xi_{m}^{*}\left(\omega\right)\xi_{n}\left(\omega\right)\mathrm{d}\omega,$ (7)
    where $pm,qn$ corresponds to the matrix rows and columns according to feeding lines $p,q$ and basis function $m,n$, respectively. Kronecker delta $\delta_{pq}$ is zero except when $p$ and $q$ refer to the same port. Since the ports are separated and the basis functions are orthogonal to each other with respect to the given integral, the identity matrix represents the total incident energy matrix $\mathbf{W}$. Equivalently projction of other operators considered in this methodological example are presented in Appendix C.
@@
    The optimization task is resolved using standard methods of the QCQP. The optimization problem is convex, so the solution is unique [14], which ensures that the maximum is truly achieved [^5] The optimization problem is equivalent to that for determining performance limitations on antenna metrics in the frequency domain [19].. Within this paper, the optimization is more than a thousand times faster than the computation of the impulse response of the system. For more details about the convex optimization framework, see [39].
@@
Although described for far-field radiation scenarios, the methodology explained above is general and can easily be applied to different directions and extended to applications such as near-field energy focusing, wireless power transfer, or UWB radars. However, we intentionally omit design concerns related to these systems, such as topology optimization of the radiating and receiving antennas. We focus solely on feeding waveform optimization, which is, unlike topology optimization, a convex optimization problem.
@@ III-B Verification
The steps mentioned above are applied to an example of a single dipole antenna. The calculation is simultaneously performed in the Antenna Toolbox for MATLAB [40] (AToM) and CST. The dipole with a length of $15\,\mathrm{mm}$ and a width of $3\,\mathrm{mm}$ is oriented along the $z$ axis. The port is placed in the middle of the dipole and is connected to an ideal $50\,\mathrm{\Omega}$ feed line. Radiation in the spherical coordinate system is studied in the direction $\varphi=0$ and $\theta=\pi/2$. Azimuthal coordinate $\varphi$ references to angle from positive direction of $x$ axis in $xy$ plane and elevation coordinate $\theta$ references to angle from positive direction of $z$ axis.
The antenna, including the port, is given, and all steps visualized in Fig. 1 can be performed. Choosing AToM and CST as the EM solvers and computing the far-field component $F_{\theta}$ in elevation direction for unity excitation at all frequencies results in Fig. 2.
@@ IV-A1 Singleport Antennas
In [43], the conical spiral antenna is proposed for ground-penetrating radars, an UWB application. The conical spiral radiates a circularly polarized field compared to the other designs which operate in a linearly polarized regime. The antenna proposed in [43] is scaled to fit the circumscribing sphere mentioned above. The simulation is again performed in CST and shown in Fig. 7.
@@
The periodic logarithmic antenna, recognized as an UWB antenna [44], is extensively utilized in electromagnetic emission and immunity testing [6]. Therefore, a wire dipole array is a potential candidate for pulsed radiation and is compared with other antennas. The array is designed to fit the circumscribing sphere and exhibits a decay of $0.9$ in element lengths and an opening angle of $15$ degrees. The strip width equals $0.017$ times the length of the longest element and the number of elements is 20. The antenna is simulated using AToM and shown in Fig. 8.
@@
The comparison of the antennas and the finding of the optimal excitation using the approach proposed in this paper show that antenna design and pulse generation capabilities are closely related. The information that can be provided by this computational approach can provide essential insight into designing antennas that perform well in pulsed radiation.
Notably, within the optimization framework (1), constraints on field intensity at spatial points can be included without compromising generality or convexity. Similarly, upcoming sections demonstrate the imposition of energy constraints at the target point within a time window. This method can also be extended to additional spatial points outside the main direction and time frames, still using the same framework. However, the inclusion of these constraints may result in an increase in the duration of the optimization process.
@@ IV-A2 Multiport Systems
The extension to multiport antennas or arrays is straightforward, as with dipole arrays within this section. The matrix formulation is also shown in (7), Appendices A and C.
@@
Considering the same input energy, the end-fire array increases the performance by more than a factor of four as compared to a single dipole. In the case of a broadside array, the performance boost reaches a factor of six. The difference may be due to the coupling of the field with the elements of the arrays which can be deduced from the optimal excitation waveforms in Fig. 13. Although the end-fire and broadside arrays share the same configuration, their characteristics, especially impedance ratios, differ since the dipoles are fed simultaneously. In this particular case, the broadside direction performs better than the end-fire orientation.
@@
The suggested approach can be applied to multi-dimensional arrays or any radiating structures with multiple ports and arbitrary configurations. Typically, computational demands grow as the number of elements increases. As demonstrated, expanding the number of independent ports results in expressing the excitation waveforms at each port through the basis functions described in Appendix B. The expansion coefficients are distinctively grouped into a vector that includes all dof.
@@ IV-B Switching Antiferromagnetic Memory
Figure 15: Incident plane waves at the position of the observation: (o) Pulse used in [45]. (a) Optimal excitation for the maximum peak. (b) Optimal excitation for the maximum peak with a constraint on the incident wave to concentrate $90\,\%$ of energy in the $2~{}\mathrm{ps}$ interval. (c) Optimal excitation for the maximum peak with a constraint in the target field to concentrate $90\,\%$ of energy in the $14~{}\mathrm{ps}$ interval. (d) Optimal excitation for the maximum peak with both constraints in (b) and (c).
@@
Figure 16: Magnitude of the electric field in CuMnAs for excitation pulses (o)–(d) in Fig. 15.
@@
One of the optimal pulses presented corresponds to the maximum magnitude of the electric field in CuMnAs in $t=0$ which is denoted as (a). The optimization problem is identical to (8) other than matrix $\mathbf{S}$ being replaced by $\mathbf{E}^{\mathrm{H}}\mathbf{E}$, where the electric field at the target point is represented as $\boldsymbol{E}_{\mathrm{CuMnAs}}(\omega)=\mathbf{E}(\omega)\mathbf{q}$ in the frequency domain, and $\mathbf{q}$ corresponds to the expansion of incident plan wave $a(\omega)$ into the basis functions of Appendix B, as done in (5). The relation is computed using CST which allows us to compute the electric field at a given point as a function of frequency considering unity excitation. Mapping the set of the given basis functions results in the desired matrix. The matrix representing the total energy of the incident plane wave is derived from (4) as demonstrated in (7), with the only variation being the inclusion of the free space impedance in the denominator.
@@
where $T=1\,\mathrm{ps}$ is the half-time interval and the asterisk represents a complex conjugation. In the frequency domain, the sinc function represents the process of time windowing. The matrix obtained is the outcome of the procedure illustrated in (7) regarding total input power. The corresponding QCQP reads
@@
The combination of both constraints for excitation in (b) and for the target field in (c) results in another optimization problem whose solution is presented in (d). Since optimization has more constraints than the previous solutions (a), (b), and (c), performance is lower but still higher than that of the original pulse (o). That shows that optimization has achieved maximum potential in the given scenario considering the constraints. The excitation mostly spans the narrow interval and the electric field in CuMnAs is also mainly accumulated in the given interval. The disparity between (o) and (d) is minimal, so the used design performs almost the best possible. This suggests that enhancement can only be achieved through topological changes even more by co-design of the structure and of the pulse.
@@ V Conclusion
The tool can serve and provide quantitative results in many scenarios, including pulsed radiation or antiferromagnetic memory switching with pulse lasers, as discussed in this paper, as well as the creation of hotspots using near-field energy focusing in lossy materials using antenna arrays for applications such as hyperthermia therapy in medicine.
@@
An important strength of the proposed method lies in its practical utility for evaluating and optimizing radiating systems. Given a specific system configuration, the method determines whether the desired pulsed performance can be achieved. If the shape and constraints of the system are sufficient, the method identifies the optimal pulse shape to maximize performance. Conversely, if the system cannot meet the imposed requirements, the approach highlights these limitations, offering valuable insights into whether a redesign of the system geometry is necessary.
By decoupling waveform optimization from structural design, the proposed method ensures a globally optimal solution that can be computed efficiently in polynomial time, unlike the iterative and often local solutions associated with topology optimization. This characteristic guarantees the general applicability of the method to existing antenna systems without necessitating their redesign, while still ensuring the best achievable performance for a given scenario and set of constraints.
While this paper focuses on waveform optimization for a fixed transmitting antenna, the methodology has broader potential. It can be extended to co-design scenarios in which both antenna and waveform design are considered simultaneously, making it applicable to a wider range of advanced engineering problems. In such case, the topology optimization would change the design according to the results computed by the proposed tool, though such a procedure would be computationally intensive. However, this would open the way to the automated design of pulsed radiators, antiferromagnetic memories, or other devices that rely on pulses in the time domain.
@@ Appendix A Matrix Operators from EFIE and MoM
A representative example of an EM solver capable of providing required antenna characteristics is AToM based on EFIE and MoM. The matrix representation of multiple quantities is presented in [39, Appendix A]. The extraction of essential metrics to the manuscript is presented in this appendix. Multiple ports are considered.
@@
$\boldsymbol{E}_{\infty}(\boldsymbol{r},\omega)=\mathbf{F}(\boldsymbol{r},\omega)\mathbf{I}(\omega)\dfrac{\mathrm{e}^{-\mathrm{j}\omega r/c}}{r},$ (13)